### v2

# Growth and characterization of the La3Ni2O7-δ thin films: dominant contribution of the orbital at ambient pressure

Abstract
By using the pulsed-laser-ablation technique, we have successfully grown the La3Ni2O7-δ thin films with -axis orientation perpendicular to the film surface. X-ray diffraction shows that the (00l) peaks can be well indexed to the La3Ni2O7-δ phase. Resistive measurements show that the samples can be tuned from weak insulating to metallic behavior through adjusting the growth conditions. Surprisingly, all curves of in the temperature region of 2300 K do not show the anomalies corresponding to either the spin density wave or the charge density wave orders as seen in bulk samples. Hall effect measurements show a linear field dependence with the dominant hole charge carriers, but the Hall coefficient exhibits strong temperature dependence. The magnetoresistance above about 50 K is positive but very weak, indicating a weakened or absence of multiband effect. However, a negative magnetoresistance is observed at low temperatures, which shows the delocalization effect by magnetic field. Detailed analysis on the magnetoresistance suggests that the delocalization effect at low temperatures is due to the Kondo-like effect, rather than the Anderson weak localization. Our transport results suggest that, the electronic conduction is fulfilled by the orbital with holes as the dominant charge carriers, while the interaction through Hund’s coupling with the localized orbital plays an important role in the charge dynamics.
I INTRODUCTION
Since the discovery of cuprate superconductors, much efforts have been made in order to find other superconductors which are analogous to the cuprates. Ni is adjacent to Cu in the periodic table, and some of Ni-based compounds share many similarities with cuprate superconductors in terms of crystal structure and band structures. Thus, researchers have pointed out that high-temperature superconductivity may be expected in nickelates [1, 2, 3, 4]. However, many trial experiments along this line did not lead to success. Until 2019, superconductivity with a transition temperature () around 915 K was discovered in infinite-layer Nd1-xSrxNiO2 thin films [5] in which the Ni ions are supposed to have a valence state close to +1, leading to a similar outer-shell electronic state of as in cuprates. Since then, researchers have found superconductivity in infinite-layer nickelate thin films with variety of rare earth elements, such as La, Pr, etc., and hole doping through alkaline-earth elements [6, 7, 8]. Signature of superconductivity was also found in Nd6Ni5O12 thin films in which the Ni ions have similar electron filling state as in Nd1-xSrxNiO2 thin films [9]. As far as we know, superconductivity is still absent in bulk samples of infinite-layer nickelates [10, 11]. In thin films, the highest of infinite-layer nickelates has reached 20 K at ambient pressure [12], and around 30 K at a high pressure of 12.1 GPa [13] and rises with pressure without a saturation yet.
Very recently, signature of superconductivity with around 80 K was found in bilayer nickelates La3Ni2O7 single crystals under high pressure [14], which has attracted great attention in the community. Subsequently, superconducting state with zero resistance in La3Ni2O7 under high pressure is confirmed [15]. This means that nickelates are likely to become the third type of unconventional high-temperature superconductive family after cuprates and iron-based superconductors. This discovery can also stimulate the research on the pairing mechanism of unconventional high-temperature superconductivity. Unlike cuprates and infinite-layer nickelates, the Ni ions in La3Ni2O7 exhibit a electron configuration. Theoretical calculations show that strong inter-layer coupling of orbital forms the inter-layer -bonding and anti-bonding bands which locate below and above the Fermi level, respectively. It is supposed that the bonding band crosses the Fermi level under high pressure, and the inter-layer coupling may play a crucial role in the pairing for superconductivity in La3Ni2O7 [14, 16].
Through continuous efforts of studies, superconductivity has also been observed in polycrystalline La3Ni2O7 [17], La3-xPrxNi2O7 [18] and trilayer nickelates La4Ni3O10 [19, 20, 21, 22, 23]. However, superconductivity of nickelates with these Ruddlesden-Popper (RP) structure all needs a high pressure. Recently, the phase diagram of La3Ni2O7 under pressure has been modified to a triangular shape [24]. It is expected that the minimum pressure to induce superconductivity in RP nickelates could be reduced so that more in-depth research can be conducted. A feasible solution would be the preparation and research of high-quality thin films. In thin films, the biaxial strain provided by the substrate and the influence of the interface can change not only the lattice parameters of the sample, but also the oxygen octahedral torsion in the perovskite structures [25, 26, 27, 28, 29]. A recent research indicated that the phase in RP nickelate thin films can be tuned by the strain of the substrate [30].
In this paper, we report the successful fabrication of a series of thin films with La3Ni2O7-δ as the dominant phase by pulsed laser deposition (PLD). Through adjusting the growth conditions, weak insulating and metallic behavior were obtained in the La3Ni2O7-δ thin films. Unlike the previous study that the La3Ni2O7 film with high crystalline exhibits insulating temperature dependence [31], our thin films show a clear metallic behavior. The X-ray diffraction (XRD) pattern and rocking curves of our La3Ni2O7-δ films indicate that they have good crystallinity, which may be one of the reasons for the good conduction and metallic behavior. The curves of all the samples do not show any anomalies corresponding to either the spin density wave (SDW) or the charged density wave (CDW) as seen in bulk samples [17, 15, 32, 33, 34, 35, 36, 37, 36]. We have also measured the Hall resistance and magnetoresistance (MR) of a thin film showing metallic behavior in a wide temperature region. The transverse resistivity shows linear field dependence with dominant hole charge carriers, but the Hall coefficient exhibits strong temperature dependence. The MR above 50 K is positive but very weak, and negative at low temperatures. Our analysis suggests that holes originated from the orbital are the dominant charge carriers.
II EXPERIMENTAL DETAILS
The La3Ni2O7-δ thin films were grown on single-crystal (00)-oriented LaAlO3 substrates by the technique of pulsed-laser deposition, with substrate temperature at and laser frequency of 2 Hz for all cases. Prior to the deposition, a polycrystalline La3Ni2O7-δ target was prepared by the sol-gel method [38]. During the growth, the laser spot area of 4.68 was fixed, and two laser fluence of 1.52 (labelled as SL) and 2.07 (labelled as SH) were applied. And the oxygen partial pressure during the growth was 2035 Pa. After growth, the samples were cooled down at a cooling rate of to under oxygen partial pressure =58 Pa and post-annealed in-situ at for about 148 h, followed by the same cooling procedure to room temperature. Then they were taken out for the subsequent measurements. XRD and X-ray reflectivity (XRR) measurements were performed by using a Bruker D8 Advanced diffractometer with Cu radiation. The reciprocal space map (RSM) is measured by a Bruker D8 discover diffractometer. The morphology characterization obtained by a scanning electron microscope (SEM) (Fig. 2(a)) was conducted by using the secondary electron imaging at an accelerating voltage of 10 kV. The size of the crystalline and surface roughness were obtained by an atomic force microscopy (AFM). The surface diffraction pattern was obtained by an in-situ high-pressure reflection high-energy electron diffraction (HP-RHEED) with a voltage of 30 kV after growth. Transport properties were measured by a physical property measurement system (PPMS, Quantum Design) with a magnetic field up to 9 T.
III RESULTS AND DISCUSSION
III.1 Structural characterizations
Fig. 1(a) shows the wide-range XRD 2- scans for the metallic La3Ni2O7-δ thin films (sample SH2 and SH3). All the (00) ( is even) peaks of La3Ni2O7-δ appeared without any obvious impurity peaks, suggesting the -axis orientation of the films without any visible impurity phases. Fig. 1(b)-(e) show the enlarged XRD 2- scans of as-grown films around the La3Ni2O7-δ (006) peak (b and c) and (0010) peak (d and e). The dashed lines in grey, light red, and light green represent the peak positions of , , and La4Ni3O10 phases, respectively. The XRD 2- measurements indicate that all of the thin films display the dominant La3Ni2O7-δ phase. Samples in Fig. 1(b) and (d) were grown with lower laser energy (laser fluence 1.52 ), and samples in Fig. 1(c) and (e) were grown with higher laser energy (laser fluence 2.07 ). The remarks in Fig. 1(b) and (c) mark the oxygen partial pressure for the growth of each sample. Sample SL4 and SH3 were post-annealed under oxygen partial pressure = Pa at for 24 and 48 h, respectively. Other samples were post-annealed under oxygen partial pressure = Pa at for 1 h. As shown in Fig. 1(f), the films grown with higher laser energy have better crystallinity than those with lower laser energy. Fig. 1(g) shows the X-ray reflectivity of sample SH3 with corresponding fit shown in a red dashed line, which yields a film thickness of about 12.056 nm. The thicknesses of samples SH1 and SH2 are around 20 nm. Other samples have also thickness around 12nm. Fig. 1(h) shows the reciprocal space map of sample SH3 around the (103) diffraction peak of the LaAlO3 substrate. The dashed line indicates that the La3Ni2O7-δ thin film has a in-plane constrain by the substrate. The diffraction peak broadening on the left-hand side of the La3Ni2O7-δ thin film is slightly stronger than that on the right-hand side, which may be due to a slight relaxation in the surface layer. The -axis constant is calculated to be between 3.792 Å and 3.807 Å, smaller than that of the bulk samples. And the -axis constant is around 20.722 Å, slightly larger than that of the bulk sample due to the in-plain compressive strain.
Fig. 2(a) shows the scanning electron microscopy secondary electron image of sample SH3. The surface of the film is relatively flat. Fig. 2(b) shows the atomic force microscopy topography of sample SH3. The diameter of the grains is about 50100 nm as shown in the inset of Fig. 2(b), and the average roughness of the surface is about 0.32 nm. Fig. 2(c) shows the RHEED pattern of sample SH3. The clear diffraction streaks at high oxygen partial pressure of 35 Pa indicate a flat surface of the film.
III.2 Transport measurements
Fig. 3(a) and (b) show the curves of La3Ni2O7-δ thin films grown with lower and higher laser energy, respectively. It is clear that the resistivity of the samples decreases with the increasing oxygen partial pressure for both lower and higher deposition energy. However, as shown in Fig. 3(a), samples synthesized with lower laser energy are weak insulators for all deposition oxygen partial pressures and annealing conditions. We also tried growth with higher growth partial pressures of oxygen (40 Pa) and with lower laser fluence. But the dominant phase of the thin film becomes La4Ni3O10 instead of the metallic La3Ni2O7-δ. This may be the reason for the poor crystallinity of the La3Ni2O7-δ in the these samples, as shown in Fig. 1(f). As shown in Fig. 3(b), the samples synthesized with higher laser energy are insulating when the growth oxygen partial pressure is less than 30 Pa. With the oxygen partial pressure raised to 35 Pa, the La3Ni2O7-δ thin films demonstrate a clear metal-to-insulator transition as temperature decreases. Through long-time annealing at high oxygen partial pressure, the metal-to-insulator transition temperature can be lowered from 51.3 K (sample SH2) to 24.3 K (sample SH3). This may suggests that the weak insulating behavior of La3Ni2O7-δ is related to oxygen vacancies, consistent with previous studies [39, 40, 41, 42]. It is noteworthy that all the curves in the whole measurement temperature region do not show the DW-like or SDW-like anomalies as seen in bulk samples. The possible transition may be suppressed by the substrate strain, as that in the bulks is suppressed by physical pressure [15, 17, 32]. The inset of Fig. 3(b) shows the resistivity behavior of sample SH2 and SH3 at low temperatures on semi-log plots. A clear logarithmic increase of the resistivity with decreasing temperature is observed. This may correlate with Kondo effect or weak localization in a Quasi-two-dimensional (2D) system, as will be addressed below.
We also performed some ex-situ oxygen annealing on these samples. We annealed these samples in an atmosphere of oxygen using a thermal gravimetric analyzer. However, the samples became more insulating. This may be due to the heat-treatments which worsens the film crystallinity. We then tried to process the insulating samples at a high oxygen pressure of 8 MPa, but even after annealing for a long time (60 h) at such a high oxygen pressure, they still could not become metallic, only a structural phase transition appeared at around 250 K. Unlike the bulk samples, where oxygen treatment can change the samples from insulating to metallic, for the film sample, it depends more on the growth process rather than on the oxygen treatment concerning whether it is metallic or insulating . Compared with the rocking curves in Fig. 1(f), the full-width-at-half-maximum on the La3Ni2O7-δ (006) peaks of samples grown at lower laser energy are generally wider than those grown at higher energy. And the rocking curve peaks of metallic samples are much sharper than that of the weak insulating samples, indicating the crystallinity might be important to gain metallic La3Ni2O7-δ thin films.
As shown in the inset of Fig. 4(a), the raw data of the transverse resistivity shows good linear field dependence with a positive slope at all temperatures, indicating that the hole-like charge carriers dominate the electric conduction. The band structure calculation shows that the hole pockets are mainly contributed by orbital [14, 43, 44, 45]. These results suggest holes originating from orbital are the dominant charge carriers. Judged from the data in Fig. 4(a), the value of Hall coefficient is comparable to that of the single crystal at pressures below 5 GPa [46]. It is known that for a normal metal with Fermi liquid feature, the Hall coefficient should be a constant versus temperature. However, the of the La3Ni2O7-δ films exhibits strong temperature dependence that decreases with the increasing temperature, this may be related to multiband effect or correlation effect. We did the calculation of the Hall angle that , which measures only the scattering rate and the information about the charge-carrier density is naturally separated away. Here the is the scattering time determined from Hall angle, is the circling frequency. The temperature dependence of is shown in Fig. 4(b). And it can be well fitted by the expression of with . The temperature dependence of calculated from resistivity is also shown in Fig. 4(b). exhibits a complex structure, which can reflect the combined results of the true scattering rate and the temperature dependent density of states (DOS). We also try to fit the metallic part of with , and a good fitting result with is obtained. The discrepancy between and manifests that the effective DOS may have a strong temperature dependence that is depleted in lowering temperature, which is a common feature in correlated materials. And this might be part of the reason that the Hall coefficient exhibits strong temperature dependence.
To further investigate the electronic scattering and the information of the Fermi surface, we systematically measured the magnetoresistance of the metallic La3Ni2O7-δ thin film. The magnetoresistivity is defined as , where is the longitudinal resistivity in a magnetic field and is that at a zero-field. As shown in Fig. 5(a), the magnetoresistance above 50 K is positive but very weak (less than 0.1), indicating a weakened or absence of multiband effect. Thus, we think that the strong temperature dependence of the Hall coefficient should be partially contributed by the multiband effect which becomes weaker in high temperature region (due to temperature enhanced interband scattering), and partially due to the correlation effect, because the effective DOS at Fermi energy is temperature-dependent and becomes depleted with decreasing temperature. The negative magnetoresistance at temperatures below 40 K in Fig. 5(a) indicates the delocalization effect in a magnetic field, and it decreases with increasing temperature. The MR above 60 K is positive, and with the increasing temperature, it increases in the temperature region of 60140 K and decreases slightly at temperatures above 160 K. Fig. 5(b) shows the temperature dependence of the magnetoresistance at 9 T. It can be observed that the negative magnetoresistance at low temperatures exhibits a rapid decay with increasing temperature, while the positive magnetoresistance at high temperatures is notably weak and exhibits minimal variation with temperature. The inset of Fig. 5(b) shows the curve at 0 T and 9 T. The magnetic field exerts a slight suppression on the upturn of resistivity at low temperatures in a magnetic field of 9 T. According to Kohler’s law [47], if there is only one isotropic scattering time in the transport properties of a system, the magnetoresistance , where and represent some unknown functions (assuming that the scattering rate ). Fig. 5(c) shows the positive MR plotted by following Kohler’s law, One can see that Kohler’s law is obeyed at temperatures above about 180 K, disobeyed at temperatures below 160 K. As mentioned before, the variation of DOS with temperature due to correlation effect can make the basic requirement of Kohler’s law that invalid, which may cause the failure of Kohler’s law for describing the magnetoresistance.
III.3 Analysis for the resistivity and negative magnetoresistance at low temperatures
Now we turn to the negative magnetoresistance at low temperatures. It is found that this negative magnetoresistance is closely linked with the upturn of resistivity at low temperatures. The negative MR or the upturn of low-T resistivity may have two different origins: (1) Kondo effect or (2) Anderson localization in 2D systems. We first assume that the cause of the logarithmic increase in resistivity and negative MR at low temperatures is due to the Kondo effect, thus the curve at low temperatures can be fitted by
Where is the residual resistivity, is the Kondo scattering term that contributes to the logarithmic rise behavior at low temperatures, and it can be expressed by Hamann formula [48]
Where is a temperature-independent constant, is the Kondo temperature, is the spin of the magnetic impurities. In order to improve the reliability of fitting results, we first fit the curve in the temperature region of 5080 K with (since the negative magnetoresistance disappears at above 50 K and the temperature region for the Kondo fit is chosen for temperatures between 2 to 80 K), then put the value of the resulting into the Kondo fitting formula and make sure that the values of and obtained from the Kondo fit are close to those obtained from the fit for 5080 K. The curve fitted by eq. 2 of the low-temperature resistivity is shown as the green solid line in Fig. 6, with K and .
As shown by the red solid line in Fig. 6, it gives K and , being close to the fitting results by eq. 2. Since it is a fit with multiple fitting parameters, the final fitting result may subject to some errors, but the Kondo-like scattering picture seems to be substantiated by the general behavior of the integrated data system.
In a Kondo system, for , the leading term of the magnetoresistance can be given by the square of the magnetization to both second [51] and third order [52] perturbation expansions of the - exchange Hamiltonian. Considering that the magnetization can be described by the Brillouin function, we have
Where is a constant, is the Brillouin function, is the magnetic flux density, is the magnetic impurity spin, is the Lander factor, is the Bohr magneton and is the Boltzmann constant. Here we assume that the orbital angular momentum is completely frozen since the influence of the crystalline field. It should be noted that the applicability of eq. 4 is limited to temperature , since at temperatures below , the magnetic moment of the impurity ions in the metal is gradually shielded and canceled out by the electron spin [53].
Fig. 7 shows as a function of . The curves for K overlap each other. This means that the Kondo temperature is between 67.5 K, slightly smaller than the value obtained from the fits of the curve. The lines in Fig. 7 represent the Brillouin function fits (eq. 4) for K. The raw data for K lies between the fitting curves of the Brillouin function with and , which means that is likely to be bewteen 0.5 and 1, slightly larger than that obtained from fits of the curve. The difference is probably mainly due to the large errors of the fits for the curve, since there are too many parameters. As metioned in the previous references [16, 42, 54, 55, 56, 57], the Ni electrons are localized forming the localized magnetic moment, the Ni electrons are itinerant. The interaction between electrons of these two orbitals may give rise to the Kondo effect. On the other hand, the fit of the magnetoresistance at low temperatures support that the spin of the magnetic impurities is close to 1. Since Ni is the only magnetic element in the La3Ni2O7-δ thin film and its divalent ion has a spin , there may be a small amount of impurities in the sample, causing the Kondo effect.
As mentioned above, Anderson weak localization effect may also cause a logarithmic increase of resistivity and negative MR at low temperatures. In the case for a 2D weak localization, we have [58]
The inelastic scattering length is defined as , where depends on the scattering mechanism. For the major inelastic scattering due to electron-phonon scattering, we have ; for electron-electron scattering, we have [58]. is the Drude conductivity, where the term represents the contribution of Fermi liquid behavior. represents the lower length cutoff for diffusive motion and is related to the transport mean free path. Fig. 8(a) shows the sheet conductance versus ln() of the metallic La3Ni2O7-δ thin films at low temperatures, and the solid lines show the fits with eq. 5. For sample SH3, we have , but for sample SH2, , which does not make sense. Considering that the mechanism that causes the resistivity upturn shouldn’t be different for these two similar samples, we exclude the weak localization effect in the La3Ni2O7-δ films.
Furthermore, weak localization effects generally appear in two-dimensional electronic gases or metallic thin films. The former is strictly a two-dimensional system, the latter can be regarded as two-dimensional when the film thickness is less than inelastic diffusion length. When a magnetic field perpendicular to the plane is applied, due to the change of electron relative phase, the intensity of weak localization decreases rapidly. Conversely, when the applied magnetic field ia parallel to the plane, the intensity of weak localization will not vary greatly. Unlike the weak localization effect, the magnetic impurities are randomly distributed in the crystal in Kondo system. Therefore, the variation of magnetoresistance in different magnetic field directions should be similar for the Kondo-like scattering. Fig. 8(b) shows the magnetoresistance for different magnetic field directions. In the two measured directions, the negative MR has similar variation tendencies, with only a slight difference in quantity. This also supports the Kondo-like explanation.
IV SUMMARY
In summary, we have successfully synthesized (00)-oriented La3Ni2O7-δ thin films on LaAlO3 single crystal substrates by PLD technique. We have obtained samples with weak insulating and metallic behavior by adjusting the laser energy and oxygen partial pressure during growth. The rocking curves show that the crystallinity of the metallic samples is significantly better than that of the weak insulating samples. Moreover, the curves of all the La3Ni2O7-δ thin films in the temperature region of 2300 K do not show any anomalies corresponding to the spin or the charge density waves that have been observed in bulks. This may be attributed to the influence of the compressive strain provided by the substrate. Hall resistance shows a linear field dependence with the dominant hole charge carriers, but the Hall coefficient exhibits strong temperature dependence. The MR above about 50 K is positive but very weak, indicating a weakened or absence of multiband effect. The Hall angle shows that the scattering rate is close to , but the scattering rate derived from resistivity shows the temperature dependence with an exponent of about 1.5, which is explained by the involvement of correlation effect in the La3Ni2O7-δ thin films. Moreover, the magnetoresistance at low temperatures is negative, showing the delocalization effect. Detailed analysis of the magnetoresistance suggests that the delocalization effect can be attributed to the Kondo-like scattering, rather than the Anderson weak localization. The Kondo-like scattering suggests a Hund’s coupling between the orbital with the orbital.
Acknowledgements.
We thank B. Yang and Y. Z. Tian for assistance with XRD and X-ray reflectivity analysis. Thank B. Hao for help with the RSM measurement and analysis. We acknowledge financial support from the National Key Research and Development Program of China (No. 2022YFA1403201), National Natural Science Foundation of China (Nos. 11927809, 12061131001).References
- Anisimov et al. [1999] V. Anisimov, D. Bukhvalov, and T. Rice, Phys. Rev. B 59, 7901 (1999).
- Lee and Pickett [2004] K.-W. Lee and W. Pickett, Phys. Rev. B 70, 165109 (2004).
- Poltavets et al. [2010] V. V. Poltavets, K. A. Lokshin, A. H. Nevidomskyy, M. Croft, T. A. Tyson, J. Hadermann, G. Van Tendeloo, T. Egami, G. Kotliar, N. ApRoberts-Warren, et al., Phys. Rev. Lett 104, 206403 (2010).
- Zhang et al. [2017] J. Zhang, A. Botana, J. Freeland, D. Phelan, H. Zheng, V. Pardo, M. Norman, and J. Mitchell, Nat. Phys. 13, 864 (2017).
- Li et al. [2019] D. Li, K. Lee, B. Y. Wang, M. Osada, S. Crossley, H. R. Lee, Y. Cui, Y. Hikita, and H. Y. Hwang, Nature 572, 624 (2019).
- Osada et al. [2020] M. Osada, B. Y. Wang, B. H. Goodge, K. Lee, H. Yoon, K. Sakuma, D. Li, M. Miura, L. F. Kourkoutis, and H. Y. Hwang, Nano Lett 20, 5735 (2020).
- Osada et al. [2021] M. Osada, B. Y. Wang, B. H. Goodge, S. P. Harvey, K. Lee, D. Li, L. F. Kourkoutis, and H. Y. Hwang, Adv. Mater 33, 2104083 (2021).
- Zeng et al. [2022] S. Zeng, C. Li, L. E. Chow, Y. Cao, Z. Zhang, C. S. Tang, X. Yin, Z. S. Lim, J. Hu, P. Yang, et al., Sci. Adv 8, eabl9927 (2022).
- Pan et al. [2022] G. A. Pan, D. Ferenc Segedin, H. LaBollita, Q. Song, E. M. Nica, B. H. Goodge, A. T. Pierce, S. Doyle, S. Novakov, D. Córdova Carrizales, et al., Nat. Mater 21, 160 (2022).
- Li et al. [2020a] Q. Li, C. He, J. Si, X. Zhu, Y. Zhang, and H.-H. Wen, Commun. Mater. 1, 16 (2020a).
- Wang et al. [2020] B.-X. Wang, H. Zheng, E. Krivyakina, O. Chmaissem, P. P. Lopes, J. W. Lynn, L. C. Gallington, Y. Ren, S. Rosenkranz, J. Mitchell, et al., Phys. Rev. Mater 4, 084409 (2020).
- Sun et al. [2023a] W. Sun, Y. Li, R. Liu, J. Yang, J. Li, W. Wei, G. Jin, S. Yan, H. Sun, W. Guo, et al., Adv. Mater 35, 2303400 (2023a).
- Wang et al. [2022] N. Wang, M. Yang, Z. Yang, K. Chen, H. Zhang, Q. Zhang, Z. Zhu, Y. Uwatoko, L. Gu, X. Dong, et al., Nat. Commun 13, 4367 (2022).
- Sun et al. [2023b] H. Sun, M. Huo, X. Hu, J. Li, Z. Liu, Y. Han, L. Tang, Z. Mao, P. Yang, B. Wang, et al., Nature 621, 493 (2023b).
- Zhang et al. [2024a] Y. Zhang, D. Su, Y. Huang, Z. Shan, H. Sun, M. Huo, K. Ye, J. Zhang, Z. Yang, Y. Xu, et al., Nat. Phys pp. 1–5 (2024a).
- Lu et al. [2024] C. Lu, Z. Pan, F. Yang, and C. Wu, Phys. Rev. Lett 132, 146002 (2024).
- Wang et al. [2024] G. Wang, N. Wang, X. Shen, J. Hou, L. Ma, L. Shi, Z. Ren, Y. Gu, H. Ma, P. Yang, et al., Phys. Rev. X 14, 011040 (2024).
- Wang et al. [2023] G. Wang, N. Wang, Y. Wang, L. Shi, X. Shen, J. Hou, H. Ma, P. Yang, Z. Liu, H. Zhang, et al. (2023), eprint arXiv:2311.08212.
- Zhu et al. [2024] Y. Zhu, E. Zhang, B. Pan, X. Chen, D. Peng, L. Chen, H. Ren, F. Liu, N. Li, Z. Xing, et al. (2024), eprint arXiv:2311.07353.
- Li et al. [2024a] Q. Li, Y.-J. Zhang, Z.-N. Xiang, Y. Zhang, X. Zhu, and H.-H. Wen, Chinese Phys. Lett 41, 017401 (2024a).
- Sakakibara et al. [2024] H. Sakakibara, M. Ochi, H. Nagata, Y. Ueki, H. Sakurai, R. Matsumoto, K. Terashima, K. Hirose, H. Ohta, M. Kato, et al., Phys. Rev. B 109, 144511 (2024).
- Zhang et al. [2024b] M. Zhang, C. Pei, X. Du, W. Hu, Y. Cao, Q. Wang, J. Wu, Y. Li, H. Liu, C. Wen, et al. (2024b), eprint arXiv:2311.07423.
- Li et al. [2024b] J. Li, C.-Q. Chen, C. Huang, Y. Han, M. Huo, X. Huang, P. Ma, Z. Qiu, J. Chen, X. Hu, et al., Sci. China: Phys. Mech. Astron. 67, 117403 (2024b).
- Li et al. [2024c] J. Li, P. Ma, H. Zhang, X. Huang, C. Huang, M. Huo, D. Hu, Z. Dong, C. He, J. Liao, et al. (2024c), eprint arXiv:2404.11369.
- Borisevich et al. [2010] A. Y. Borisevich, H. J. Chang, M. Huijben, M. P. Oxley, S. Okamoto, M. K. Niranjan, J. D. Burton, E. Y. Tsymbal, Y.-H. Chu, P. Yu, et al., Phys. Rev. Lett 105, 087204 (2010).
- Aso et al. [2013] R. Aso, D. Kan, Y. Shimakawa, and H. Kurata, Sci. Rep 3, 2214 (2013).
- Moon et al. [2014] E. Moon, P. Balachandran, B. J. Kirby, D. Keavney, R. Sichel-Tissot, C. Schleputz, E. Karapetrova, X. Cheng, J. M. Rondinelli, and S. May, Nano Lett 14, 2509 (2014).
- Fowlie et al. [2017] J. Fowlie, M. Gibert, G. Tieri, A. Gloter, J. Íñiguez, A. Filippetti, S. Catalano, S. Gariglio, A. Schober, M. Guennou, et al., Adv. Mater 29, 1605197 (2017).
- Fowlie et al. [2019] J. Fowlie, C. Lichtensteiger, M. Gibert, H. Meley, P. Willmott, and J.-M. Triscone, Nano Lett 19, 4188 (2019).
- Cui et al. [2023] T. Cui, S. Choi, T. Lin, C. Liu, G. Wang, N. Wang, S. Chen, H. Hong, D. Rong, Q. Wang, et al. (2023), eprint arXiv:2311.13228.
- Li et al. [2020b] Z. Li, W. Guo, T. Zhang, J. Song, T. Gao, Z. Gu, and Y. Nie, APL Mater 8 (2020b).
- Khasanov et al. [2024] R. Khasanov, T. J. Hicken, D. J. Gawryluk, L. P. Sorel, S. Bötzel, F. Lechermann, I. M. Eremin, H. Luetkens, and Z. Guguchia (2024), eprint arXiv:2402.10485.
- Kakoi et al. [2024] M. Kakoi, T. Oi, Y. Ohshita, M. Yashima, K. Kuroki, T. Kato, H. Takahashi, S. Ishiwata, Y. Adachi, N. Hatada, et al., J. Phys. Soc. Jpn 93, 053702 (2024).
- Liu et al. [2023] Z. Liu, H. Sun, M. Huo, X. Ma, Y. Ji, E. Yi, L. Li, H. Liu, J. Yu, Z. Zhang, et al., Sci. China: Phys. Mech. Astron. 66, 217411 (2023).
- Hosoya et al. [2008] T. Hosoya, K. Igawa, Y. Takeuchi, K. Yoshida, T. Uryu, H. Hirabayashi, and H. Takahashi, in J. Phys. Conf. Ser (IOP Publishing, 2008), vol. 121, p. 052013.
- Wu et al. [2001] G. Wu, J. Neumeier, and M. Hundley, Phys. Rev. B 63, 245120 (2001).
- Fukamachi et al. [2001] T. Fukamachi, K. Oda, Y. Kobayashi, T. Miyashita, and M. Sato, J. Phys. Soc. Jpn 70, 2757 (2001).
- Poltavets et al. [2006] V. V. Poltavets, K. A. Lokshin, T. Egami, and M. Greenblatt, Mater. Res. Bull. 41, 955 (2006).
- Kobayashi et al. [1996] Y. Kobayashi, S. Taniguchi, M. Kasai, M. Sato, T. Nishioka, and M. Kontani, J. Phys. Soc. Jpn 65, 3978 (1996).
- Ling et al. [2000] C. D. Ling, D. N. Argyriou, G. Wu, and J. Neumeier, J. Solid State Chem 152, 517 (2000).
- Zhang et al. [1994] Z. Zhang, M. Greenblatt, and J. Goodenough, J. Solid. State Chem (United States) 108 (1994).
- Taniguchi et al. [1995] S. Taniguchi, T. Nishikawa, Y. Yasui, Y. Kobayashi, J. Takeda, S.-i. Shamoto, and M. Sato, J. Phys. Soc. Jpn. 64, 1644 (1995).
- Yang et al. [2024] J. Yang, H. Sun, X. Hu, Y. Xie, T. Miao, H. Luo, H. Chen, B. Liang, W. Zhu, G. Qu, et al., Nat. Commun 15, 4373 (2024).
- Luo et al. [2023a] Z. Luo, X. Hu, M. Wang, W. Wú, and D.-X. Yao, Phys. Rev. Lett 131, 126001 (2023a).
- Luo et al. [2023b] Z. Luo, B. Lv, M. Wang, W. Wú, and D.-X. Yao (2023b), eprint arXiv:2308.16564.
- Zhou et al. [2024] Y. Zhou, J. Guo, S. Cai, H. Sun, P. Wang, J. Zhao, J. Han, X. Chen, Y. Chen, Q. Wu, et al. (2024), eprint arXiv:2311.12361.
- Ziman [2001] J. M. Ziman, Electrons and phonons: the theory of transport phenomena in solids (Oxford university press, 2001).
- Hamann [1967] D. R. Hamann, Phys. Rev. 158, 570 (1967).
- Goldhaber-Gordon et al. [1998] D. Goldhaber-Gordon, J. Göres, M. Kastner, H. Shtrikman, D. Mahalu, and U. Meirav, Phys. Rev. Lett 81, 5225 (1998).
- Lee et al. [2011] M. Lee, J. Williams, S. Zhang, C. D. Frisbie, and D. Goldhaber-Gordon, Phys. Rev. lett 107, 256601 (2011).
- Yosida [1957] K. Yosida, Phys. Rev. 107, 396 (1957).
- Béal-Monod and Weiner [1968] M.-T. Béal-Monod and R. Weiner, Phys. Rev. 170, 552 (1968).
- Wilson [1975] K. G. Wilson, Rev. Mod. Phys 47, 773 (1975).
- Qu et al. [2024] X.-Z. Qu, D.-W. Qu, J. Chen, C. Wu, F. Yang, W. Li, and G. Su, Phys. Rev. Lett 132, 036502 (2024).
- Ouyang et al. [2024] Z. Ouyang, J.-M. Wang, J.-X. Wang, R.-Q. He, L. Huang, and Z.-Y. Lu, Phys. Rev. B 109, 115114 (2024).
- Cao and Yang [2024] Y. Cao and Y.-f. Yang, Phys. Rev. B 109, L081105 (2024).
- Oh and Zhang [2023] H. Oh and Y.-H. Zhang, Phys. Rev. B 108, 174511 (2023).
- Lee and Ramakrishnan [1985] P. A. Lee and T. Ramakrishnan, Rev. Mod. Phys 57, 287 (1985).